# STIV: Scalable Text and Image Conditioned Video Generation

### v1

@@ 3.1 Basic Setup
Evaluation We mainly use VBench [31], VBench-I2V and MSRVTT [63] to evaluate T2V and TI2V models. For VBench, we mainly report Quality (temporal quality and frame-wise quality), Semantic (semantic alignment with different perspectives of the input text prompt) and Total score (weighted average of Quality and Semantic), and they can be actually decomposed into 16 dimensions in total. VBench-I2V builds upon the VBench with three new Video-Image Alignment metrics: Subject Consistency, Background Consistency, and Camera Motion Control. These additional metrics provide a more comprehensive evaluation by focusing on how well the generated video aligns with the input image and specified prompt instructions. More details about the detailed dimensions are presented in Section F.1. We present three model scales: XL, XXL, and M with their configuration detailed in Tab. 1. In the following section, we use the notation X-S-T to represent an X-size model with an $S^{2}$ resolution and T frames. If unspecified, the default configuration is a $256^{2}$ resolution with 20 frames. More detailed model and training configurations are provided in the Appendix.
@@ 3.4.2 The Effectiveness of Image Condition Dropout
As shown in the Fig. 7, we observed: (1) Fixing CFG-T, as CFG-I increases, FVD first decreases and then increases; (2) Fixing CFG-I, as CFG-T increases, FVD continuously decreases, except when CFG-I is very small (1.1, 1.5), where it first decreases and then increases; (3) The best configuration occurs at $CFG-T=7.5$ and $CFG-I=1.5$, yielding $FVD=95.2$. However, overall, SIT-CFG does not show significant advantages compared to JIT-CFG, and using two copies for inference is significantly less efficient. Note that this search was optimized for MSRVTT, and for other prompts requiring stronger dependence on the first-frame subject, a larger CFG-I might be needed.
@@ 3.5 Video Data Engine
[…]
  Caption Total Object DSG-$\text{Video}_{i}(\downarrow)$ DSG-$\text{Video}_{s}(\downarrow)$ MSRVTT FVD ($\downarrow$) VBench ($\uparrow$)
@@ 4 Results
Based on all of these studies, we scale our T2V and STIV model from 600M to 8.7B. We show the main results in Tab. 9 and Tab. 10, comparing our models with state-of-the-art open sourced and close sourced models, which demonstrates the effectiveness of our recipes. Specifically, we do finetuning on top of the pretrained video generation models (SFT), based on the 20,000 videos filtered from Panda-70M [10] using the method mentioned in Section 3.5. Since we adopt MaskDiT technique in our pretraining stage, we try finetuning our model in an unmask manner (UnmaskSFT). We also finetuned our STIV model to become a temporal upsampler to interpolate the videos generated by our main T2V and STIV models to boost the motion smoothness (+ TUP).
T2V Performance We first showcase the effectiveness of our T2V model as the foundation for STIV. Tab. 9 presents a comparison of different T2V model variants on VBench, including the VBench-Quality, VBench-Semantic, and VBench-Total scores. Our analysis reveals that scaling up model parameters in our T2V model improves semantic following capability. Specifically, as model size increase from XL to XXL and M, VBench-Semantic scores rise from 72.5 to 72.7 and then to 74.8. This explicit emergence (from XL, XXL to M), suggesting larger models are better at capturing semantic information. However, the impact on video quality, measured by VBench-Quality, remains modest, with only a slight increase from 80.7 to 81.2 and then to 82.1. This finding suggests that scaling has a greater effect on the model’s semantic capabilities than on video quality. Furthermore, increasing the spatial resolution from 256 to 512 significantly boosts the VBench-Semantic score from 74.8 to 77.0. Detailed results can be found in  Tab. 11.
@@
TI2V Performance As shown in Tab. 10, our model delivers competitive performance compared to state-of-the-art approaches. It also reveals that while scaling up improves the I2V score, it has minimal impact on quality. In contrast, increasing the resolution leads to noticeable improvements in both quality and I2V scores. We provide complete results for the decomposed dimensions in  Tab. 12.
@@ 5 Flexible Applications
Here, we adopt the novel view camera definitions outlined in Zero123++ [55], which specifies six novel view cameras for each input image. The initial frame in our TI2V model is set as the given image, and the next six frames, representing novel views, are predicted as future frames within TI2V. For training, we began with our TI2V-XL checkpoint trained with a 256 resolution, fine-tuning it for 110k steps on Objaverse [16]. For a fair comparison, we increased the image resolution to 320 during finetuning, aligning with the settings used in Zero123++. Our evaluation used objects from the Google Scanned Objects dataset [18], where we compared the output multi-view images against ground-truth renderings. As shown in Fig. 10(c), despite only using temporal attention for cross-view consistency, our approach achieves comparable performance to Zero123++ which uses full attention to all the views. This outcome validates the effectiveness of our spatiotemporal attention in maintaining 3D consistency. A visual comparison between our approach and Zero123++ is shown in Figure 11.
@@ 8 Acknowledgement
We thank Alex Schwing, Monica Zuendorf, Saeed Khorram, Pengsheng Guo and Zhe Gan for their regular discussions on model design and training recipes. We also acknowledge the indirect influence of these discussions on our approach. Furthermore, we thank Meng Cao for data and Ruoming Pang for training infrastructure support. We also acknowledge the leadership guidance from Yang Zhao throughout the project. Finally, we thank the Axlearn team  [^8] https://github.com/apple/axlearn at Apple for providing training infrastructure, which greatly facilitated our experiments.
@@ Appendix A Joint Image-Text Classifier-free Guidance
CFG-Renormalization Empirically, we observed that the magnitude of the modified velocity, $||\hat{\bm{F}}_{\theta}(\bm{x}_{t},\bm{c}_{T},\bm{c}_{I},t)||$ tends to be very large during the early stages of integration in inference (i.e. when $t$ is small). This behavior sometimes leads to overshooting beyond the learned latent distribution, resulting in artifacts in the generated output. We identified this issue as primarily due to the significant difference between the conditional velocity, $\bm{F}_{\theta}(\bm{x}_{t},\bm{c}_{T},\bm{c}_{I},t)$, and the unconditional velocity, ${\bm{F}}_{\theta}(\bm{x}_{t},\varnothing,\varnothing,t)$ when $t$ is small.
To mitigate this, we propose a simple yet effective renormalization method that re-scales the magnitude of the modified velocity to $||\bm{F}_{\theta}(\bm{x}_{t},\bm{c}_{T},\bm{c}_{I},t)||$ while preserving its direction. Formally, this is defined as:
$\tilde{\bm{F}}_{\theta}(\bm{x}_{t},\bm{c}_{T},\bm{c}_{I},t)=||\bm{F}_{\theta}(\bm{x}_{t},\bm{c}_{T},\bm{c}_{I},t)||\frac{\hat{\bm{F}}_{\theta}(\bm{x}_{t},\bm{c}_{T},\bm{c}_{I},t)}{||\hat{\bm{F}}_{\theta}(\bm{x}_{t},\bm{c}_{T},\bm{c}_{I},t)||}$
@@ Appendix I Details of Video Data Engine
[…]

### v2

@@ 3.1 Basic Setup
Evaluation We mainly use VBench [31], VBench-I2V and MSRVTT [63] to evaluate T2V and TI2V models. For VBench, we mainly report Quality (temporal quality and frame-wise quality), Semantic (semantic alignment with different perspectives of the input text prompt) and Total score (weighted average of Quality and Semantic), and they can be actually decomposed into 16 dimensions in total. VBench-I2V builds upon the VBench with three new Video-Image Alignment metrics: Subject Consistency, Background Consistency, and Camera Motion Control. These additional metrics provide a more comprehensive evaluation by focusing on how well the generated video aligns with the input image and specified prompt instructions. More details about the detailed dimensions are presented in Section F.1. We present three model scales: XL, XXL, and M with their configuration detailed in Tab.˜1. In the following section, we use the notation X-S-T to represent an X-size model with an $S^{2}$ resolution and T frames. If unspecified, the default configuration is a $256^{2}$ resolution with 20 frames. More detailed model and training configurations are provided in the Appendix.
@@ 3.4.2 The Effectiveness of Image Condition Dropout
As shown in the Fig. 7, we observed: (1) Fixing CFG-T, as CFG-I increases, FVD first decreases and then increases; (2) Fixing CFG-I, as CFG-T increases, FVD continuously decreases, except when CFG-I is very small (1.1, 1.5), where it first decreases and then increases; (3) The best configuration occurs at CFGT=7.5CFG-T=7.5 and CFGI=1.5CFG-I=1.5, yielding $FVD=95.2$. However, overall, SIT-CFG does not show significant advantages compared to JIT-CFG, and using two copies for inference is significantly less efficient. Note that this search was optimized for MSRVTT, and for other prompts requiring stronger dependence on the first-frame subject, a larger CFG-I might be needed.
@@ 3.5 Video Data Engine
Caption Evaluation and Ablations To objectively assess caption accuracy, we introduce DSG-Video(Fig. 9), a module inspired by DSG [12], that detects hallucinated objects by probing captions with LLM-generated questions and verifying object presence in sampled video frames using a multimodal LLM. This yields two metrics, $\text{DSG-Video}_{i}$ and $\text{DSG-Video}_{s}$  [^7] More details are illustrated in Appendix I., reflecting hallucination at the object and sentence levels, respectively. We compare two captioning strategies—frame-based plus LLM summarization (FCapLLM) and direct video captioning (VCap)—on the Panda-30M dataset. As shown in Tab. 8, VCap reduces hallucinations and increases the diversity of described objects, leading to improved T2V model performance. These results demonstrate that richer, more accurate video descriptions can significantly enhance downstream generation quality.
  Caption Total Object $\text{Video}_{i}(\downarrow) DSG-$$\text{Video}_{s}(\downarrow)$ MSRVTT FVD ($\downarrow$) VBench ($\uparrow$)
@@ 4 Results
Based on all of these studies, we scale our T2V and STIV model from 600M to 8.7B. We show the main results in Tab.˜9 and Tab.˜10, comparing our models with state-of-the-art open sourced and close sourced models, which demonstrates the effectiveness of our recipes. Specifically, we do finetuning on top of the pretrained video generation models (SFT), based on the 20,000 videos filtered from Panda-70M [10] using the method mentioned in Section 3.5. Since we adopt MaskDiT technique in our pretraining stage, we try finetuning our model in an unmask manner (UnmaskSFT). We also finetuned our STIV model to become a temporal upsampler to interpolate the videos generated by our main T2V and STIV models to boost the motion smoothness (+ TUP).
T2V Performance We first showcase the effectiveness of our T2V model as the foundation for STIV. Tab.˜9 presents a comparison of different T2V model variants on VBench, including the VBench-Quality, VBench-Semantic, and VBench-Total scores. Our analysis reveals that scaling up model parameters in our T2V model improves semantic following capability. Specifically, as model size increase from XL to XXL and M, VBench-Semantic scores rise from 72.5 to 72.7 and then to 74.8. This explicit emergence (from XL, XXL to M), suggesting larger models are better at capturing semantic information. However, the impact on video quality, measured by VBench-Quality, remains modest, with only a slight increase from 80.7 to 81.2 and then to 82.1. This finding suggests that scaling has a greater effect on the model’s semantic capabilities than on video quality. Furthermore, increasing the spatial resolution from 256 to 512 significantly boosts the VBench-Semantic score from 74.8 to 77.0. Detailed results can be found in  Tab.˜11.
@@
TI2V Performance As shown in Tab. 10, our model delivers competitive performance compared to state-of-the-art approaches. It also reveals that while scaling up improves the I2V score, it has minimal impact on quality. In contrast, increasing the resolution leads to noticeable improvements in both quality and I2V scores. We provide complete results for the decomposed dimensions in  Tab.˜12.
@@ 5 Flexible Applications
Here, we adopt the novel view camera definitions outlined in Zero123++ [55], which specifies six novel view cameras for each input image. The initial frame in our TI2V model is set as the given image, and the next six frames, representing novel views, are predicted as future frames within TI2V. For training, we began with our TI2V-XL checkpoint trained with a 256 resolution, fine-tuning it for 110k steps on Objaverse [16]. For a fair comparison, we increased the image resolution to 320 during finetuning, aligning with the settings used in Zero123++. Our evaluation used objects from the Google Scanned Objects dataset [18], where we compared the output multi-view images against ground-truth renderings. As shown in Fig. 10(c), despite only using temporal attention for cross-view consistency, our approach achieves comparable performance to Zero123++ which uses full attention to all the views. This outcome validates the effectiveness of our spatiotemporal attention in maintaining 3D consistency. A visual comparison between our approach and Zero123++ is shown in Figure˜11.
@@ 8 Acknowledgement
We thank Yifan Jiang,Alex Schwing, Monica Zuendorf, Saeed Khorram, Pengsheng Guo and Zhe Gan for their regular discussions on model design and training recipes. We also acknowledge the indirect influence of these discussions on our approach. Furthermore, we thank Meng Cao for data and Ruoming Pang for training infrastructure support. We also acknowledge the leadership guidance from Yang Zhao throughout the project. Finally, we thank the Axlearn team  [^8] https://github.com/apple/axlearn at Apple for providing training infrastructure, which greatly facilitated our experiments.
@@ Appendix A Joint Image-Text Classifier-free Guidance
CFG-Renormalization Empirically, we observed that the magnitude of the modified velocity, $\hat{\bm{F}}_{\theta}(\bm{x}_{t},\bm{c}_{T},\bm{c}_{I},t)||$ tends to be very large during the early stages of integration in inference (i.e. when $t$ is small). This behavior sometimes leads to overshooting beyond the learned latent distribution, resulting in artifacts in the generated output. We identified this issue as primarily due to the significant difference between the conditional velocity, $\bm{F}_{\theta}(\bm{x}_{t},\bm{c}_{T},\bm{c}_{I},t)$, and the unconditional velocity, ${\bm{F}}_{\theta}(\bm{x}_{t},\varnothing,\varnothing,t)$ when $t$ is small.
To mitigate this, we propose a simple yet effective renormalization method that re-scales the magnitude of the modified velocity to $\bm{F}_{\theta}(\bm{x}_{t},\bm{c}_{T},\bm{c}_{I},t)||$ while preserving its direction. Formally, this is defined as:
[…]
@@ Appendix I Details of Video Data Engine
Details of DSG-Video: Hallucination Evaluation To compare various captioning techniques, we develop an evaluation module that assesses both caption richness and accuracy. We quantify caption richness by measuring the diversity of unique objects referenced across captions. And we identify hallucinations to assess accuracy. Inspired by DSG [12] for fine-grained text-to-image evaluation, we introduce DSG-Video to validate the presence of objects referenced in captions within the video content (Fig. 9). First, we use an LLM to automatically generate questions that probe key details of captions, such as object identity, actions, and context. For instance, given a caption mentioning "a cat sitting on a couch", the LLM will generate questions like "Is there a cat in the video?" and "Is the cat on a couch?". Second, we utilize a multimodal LLM to answer these object verification questions by evaluating the presence of each referenced object in $N$ uniformly sampled frames from the video. For each question generated (e.g., "Is there a cat in this frame?"), the multimodal LLM examines each sampled frame and provides a response. If, for a given question, all frame-based responses suggest that the entity in question is absent, we classify the object as hallucinated. This approach ensures a thorough and frame-by-frame validation for each object in the video. Based on this, we have two metrics: one that measures the fraction of hallucinated object instances (referred to as $\text{DSG-Video}_{i}$), and the other that calculates the fraction of sentences containing hallucinated objects (referred to as $\text{DSG-Video}_{s}$). We use these two metrics to assess caption quality.
$\text{DSG-Video}_{i}=\frac{|\{\text{hallucinated objects}\}|}{|\{\text{all objects mentioned}\}|},$
$\text{DSG-Video}_{s}=\frac{|\{\text{sentences with hallucinated object}\}|}{|\{\text{ all sentences}\}|}.$